# PyramidKV: Dynamic KV Cache Compression based on Pyramidal Information Funneling

### v1

@@ 5.5 PyramidKV Excels in Memory-constrained Settings
The evaluation results from LongBench[2] are shown in Table 1 and Figure 4. In Figure 4, we report the average score across datasets for 64, 96, 128, 256 and 512 case sizes. In Table 1, we report the results for two different KV cache sizes with 128 and 2048. These two sizes represent two distinct operational scenarios—the memory-efficient scenario and the performance-preserving scenario, respectively for a trade-off between memory and model performance.
@@ 5.6 PyramidKV Preserves the Long-Context Understanding Ability
We perform Fact Retrieval Across Context Lengths (“Needle In A HayStack”) [23, 11] to test the in-context retrieval ability of LLMs after leveraging different KV cache methods. Specifically, we compare PyramidKV with SnapKV [22] using a cache size of 128 and a full KV cache size. For the medium-context cases where the input sequence is between $2,000$ and $4,000$, SnapKV produces an increasing number of bad cases in the “Needle In A HayStack” test. For the long-context cases where the input sequence is longer than $6,000$, SnapKV significantly hurts the performance of LLMs in the evaluation. In contrast, PyramidKV alleviates this weakening effect in most cases. Figure 5 shows the quantitative result. Higher scores with lighter colors indicate a better retrieval ability. The Full KV scores 65.0, PyramidKV scores 62.6, and SnapKV scores 57.3.
Figure 5: Results of the Fact Retrieval Across Context Lengths (“Needle In A HayStack”) test. The vertical axis of the table represents the depth percentage, and the horizontal axis represents the token length.  PyramidKV mitigates the negative impact of KV cache compression on the long-context understanding capability of LLMs. Higher scores indicate a better retrieval ability. For the average score across token length and depth percentage, the Full KV scores 65.0, PyramidKV scores 62.6, and SnapKV scores 57.3.

### v2

@@ 5.5 PyramidKV Excels in Memory-constrained Settings
The evaluation results from LongBench[2] are shown in Table 1 and Figure 4. In Figure 4, we report the average score across datasets for 64, 96, 128, 256 and 512 case sizes. In Table 1, we report the results for two different KV cache sizes with 128 and 2048. These two sizes represent two distinct operational scenarios—the memory-efficient scenario and the performance-preserving scenario, respectively for a trade-off between memory and model performance. In Appendix E, we report the results of KV cache sizes with 64, 96, 128, 256, 512, 1024 and 2048.
@@ 5.6 PyramidKV Preserves the Long-Context Understanding Ability
We conduct the "Fact Retrieval Across Context Lengths" (Needle In A Haystack) experiment [23, 11] to evaluate the in-context retrieval capabilities of LLMs when utilizing various KV cache compression methods. For this purpose, we employ LlaMa-3-8B-Instruct and Mistral-7B-Instruct [18] as our base, with context lengths extending up to 8k and 32k tokens. We compared several KV cache compression techniques (PyramidKV, SnapKV [22], and H2O [34]) at cache sizes of 64 and full cache. The results are displayed in Figure 5 and Figure 6 (Additional results with varying KV cache sizes with Mistral-7B-Instruct-32k-v0.2 [18] and LlaMa-3-8B-8k are available in Appendix F). The results demonstrate that with only 64 KV cache retained, PyramidKV effectively maintains the model’s ability to understand short contexts (up to a context length of 6.8k), and experiences only a modest degradation in performance on longer contexts. In contrast, other KV cache compression methods significantly impair the performance of LLMs.
Figure 5: Results of the Fact Retrieval Across Context Lengths (“Needle In A HayStack”) test in LlaMa-3-8B-Instruct with 8k context size in 64 KV cache size. The vertical axis of the table represents the depth percentage, and the horizontal axis represents the token length.  PyramidKV mitigates the negative impact of KV cache compression on the long-context understanding capability of LLMs.
Figure 6: Results of the Fact Retrieval Across Context Lengths (“Needle In A HayStack”) test in Mistral-7B-Instruct [18] with 32k context size in 64 KV cache size. The vertical axis of the table represents the depth percentage, and the horizontal axis represents the token length.  PyramidKV mitigates the negative impact of KV cache compression on the long-context understanding capability of LLMs.
@@ Appendix D License
Appendix E PyramidKV Excels in all KV Cache Size Limitation
The evaluation results from LongBench[2] are shown in Table 5. We report the results using LlaMa-3-8B-Instruct and Mistral-7B-Instruct[18] for different KV cache sizes.
Overall, consistently surpasses other method across a range of KV cache sizes and different backbone models, with its performance advantages becoming particularly pronounced in memory-constrained environments. Upon examining specific tasks, PyramidKV demonstrates a notably superior performance on the TREC task, a few-shot question answering challenge. This suggests that the model effectively aggregates information from the few-shot examples, highlighting the potential for further investigation into in-context learning tasks.
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
LlaMa-3-8B-Instruct, KV Size = Full
FKV 25.70 29.75 41.12 45.55 35.87 22.35 25.63 23.03 26.21 73.00 90.56 41.88 04.67 69.25 58.05 50.77 41.46
LlaMa-3-8B-Instruct, KV Size = 64
SKV 19.86 9.09 27.89 37.34 28.35 18.17 15.86 20.80 16.41 38.50 85.92 36.32 5.22 69.00 51.78 48.38 33.05
H2O 20.80 11.34 27.03 37.25 30.01 17.94 18.29 21.49 19.13 38.00 84.70 37.76 5.63 69.33 53.44 50.15 33.89
SLM 17.44 8.68 22.25 35.37 31.51 15.97 15.46 20.06 14.64 38.00 72.33 29.10 5.42 69.50 46.14 45.09 30.43
Ours 21.13 14.18 30.26 35.12 23.76 16.17 18.33 21.65 19.23 58.00 88.31 37.07 5.23 69.50 52.61 45.74 34.76
LlaMa-3-8B-Instruct, KV Size = 96
SKV 20.45 10.34 31.84 37.85 28.65 18.52 17.90 21.26 19.07 41.50 86.95 37.82 5.08 69.12 54.69 51.31 34.51
H2O 21.55 11.21 28.73 37.66 30.12 18.47 19.57 21.57 20.44 38.50 87.63 38.47 5.60 69.00 54.51 50.16 34.57
SLM 18.67 8.43 24.98 38.35 30.59 16.37 17.33 19.84 18.41 41.00 73.92 29.38 5.80 69.50 47.15 45.61 31.58
Ours 21.67 15.10 33.50 39.73 26.48 17.47 19.64 22.28 20.49 61.50 87.38 38.18 6.00 69.25 55.30 46.78 36.29
LlaMa-3-8B-Instruct, KV Size = 128
LlaMa-3-8B-Instruct, KV Size = 2048
Mistral-7B-Instruct, KV Size = Full
LlaMa-3-8B-Instruct, KV Size = 64
SKV 16.94 17.17 39.51 36.87 22.26 15.18 14.75 20.35 3.21 37.50 84.16 37.28 4.50 61.13 42.40 38.44 30.72
H2O 18.19 19.04 37.40 30.18 22.22 13.77 16.60 21.52 3.55 37.00 81.02 38.62 5.00 66.03 43.54 40.46 30.88
SLM 15.01 13.84 28.74 30.97 24.50 13.42 13.25 19.46 2.41 35.50 76.91 29.61 4.67 27.33 38.71 35.29 25.60
PKV 20.91 20.21 39.94 33.57 22.87 15.70 17.31 21.23 3.48 54.00 81.98 36.96 3.58 60.83 44.52 37.99 32.19
LlaMa-3-8B-Instruct, KV Size = 96
SKV 19.92 18.80 43.29 39.66 23.08 15.94 16.65 21.26 3.56 43.50 83.48 39.74 4.00 60.10 45.53 41.12 32.47
H2O 19.44 20.81 38.78 32.39 21.51 14.43 17.68 22.40 3.76 38.00 82.51 39.94 6.06 77.48 45.18 42.43 32.67
SLM 15.15 15.48 31.44 30.03 23.93 12.73 16.76 19.15 3.32 41.50 75.31 28.71 5.00 28.48 38.92 36.05 26.37
PKV 20.35 21.87 41.15 34.94 21.85 15.81 18.21 21.66 3.85 65.00 83.60 39.60 4.50 67.80 45.83 39.38 34.08
Mistral-7B-Instruct, KV Size = 128
Mistral-7B-Instruct, KV Size = 2048
SKV 25.89 32.93 48.56 42.96 27.42 19.02 26.56 24.47 4.97 70.00 86.27 42.57 5.50 88.90 50.42 46.72 40.19
Table 5: Performance comparison of  PyramidKV (Ours) with SnapKV (SKV), H2O, StreamingLLM (SLM) and FullKV (FKV) on LongBench for LlaMa-3-8B-Instruct and Mistral-7B-Instruct[18].  PyramidKV generally outperforms other KV Cache compression methods across various KV Cache sizes and LLMs. The performance strengths of  PyramidKV are more evident in small KV Cache sizes. Bold text represents the best performance.
Appendix F PyramidKV Preserves the Long-Context Understanding Ability
We perform Fact Retrieval Across Context Lengths (“Needle In A HayStack”) [23, 11] to test the in-context retrieval ability of LLMs after leveraging different KV cache methods. Figure 7, Figure 8, Figure 9 show the results of Mistral-7B-Instruct [18] with different cache size (64, 96 and 128, respectively).
Figure 10, Figure 11, Figure 12 show the results of LlaMa-3-8B with different cache size (64, 96 and 128, respectively).
Figure 7: Results of the Fact Retrieval Across Context Lengths (“Needle In A HayStack”) test in Mistral-7B-Instruct with 32k context size in 64 KV cache size. The vertical axis of the table represents the depth percentage, and the horizontal axis represents the token length.  PyramidKV mitigates the negative impact of KV cache compression on the long-context understanding capability of LLMs.
Figure 8: Results of the Fact Retrieval Across Context Lengths (“Needle In A HayStack”) test in Mistral-7B-Instruct with 32k context size in 96 KV cache size. The vertical axis of the table represents the depth percentage, and the horizontal axis represents the token length.  PyramidKV mitigates the negative impact of KV cache compression on the long-context understanding capability of LLMs.
Figure 9: Results of the Fact Retrieval Across Context Lengths (“Needle In A HayStack”) test in Mistral-7B-Instruct with 32k context size in 128 KV cache size. The vertical axis of the table represents the depth percentage, and the horizontal axis represents the token length.  PyramidKV mitigates the negative impact of KV cache compression on the long-context understanding capability of LLMs.
Figure 10: Results of the Fact Retrieval Across Context Lengths (“Needle In A HayStack”) test in LlaMa-3-8B-Instruct with 8k context size in 64 KV cache size. The vertical axis of the table represents the depth percentage, and the horizontal axis represents the token length.  PyramidKV mitigates the negative impact of KV cache compression on the long-context understanding capability of LLMs.
Figure 11: Results of the Fact Retrieval Across Context Lengths (“Needle In A HayStack”) test in LlaMa-3-8B-Instruct with 8k context size in 96 KV cache size. The vertical axis of the table represents the depth percentage, and the horizontal axis represents the token length.  PyramidKV mitigates the negative impact of KV cache compression on the long-context understanding capability of LLMs.
Figure 12: Results of the Fact Retrieval Across Context Lengths (“Needle In A HayStack”) test in LlaMa-3-8B-Instruct with 8k context size in 128 KV cache size. The vertical axis of the table represents the depth percentage, and the horizontal axis represents the token length.  PyramidKV mitigates the negative impact of KV cache compression on the long-context understanding capability of LLMs.